OpenStudy (anonymous):

Jesse’s parents bought a new car. After owning it for two years, the car’s value was $15,990. After owning it for six years, Jesse’s parents sold it to him for $2,020. Make a graph, where x is the number of years they owned the car and y is the car’s value. Graph the car’s value at 2 years and 6 years. Determine the approximate purchase price of the car.

OpenStudy (nurali):

If you assume a linear relationship between age and value. Two points: (2, 15,990) and (6, 2020) slope (m) = (y2 - y1)/(x2 - x1) = (2020 - 15,990)/(6 - 2) = -13,970/4 = -3492.5 Using m= -3492.5 and point (6, 2020): (y - y1) = m(x - x1) (y - 2020) = (-3492.5)(x - 6) y - 2020 = -3492.5x + 20,955 y = -3492.5x + 22,975 y-intercept = (0, 22,975)